OpenStudy (anonymous):

find the value of b in: f(x)=-x^2+bx-75 when the max=25

OpenStudy (anonymous):

max?

OpenStudy (anonymous):

maximum value is at 25

OpenStudy (anonymous):

try b = 20

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

well actually you can use b = 20 or b = -20

OpenStudy (anonymous):

can you show the math?

OpenStudy (anonymous):

maximum value is at the vertex. the first coordinate of the vertex is always \[-\frac{b}{2a}\] in this case since \[a=-1\] the first coordinate of the vertex is \[\frac{b}{2}\] then i replaced x by \[\frac{b}{2}\] in the equation, set the result = 25, and solved for b. would you like the steps?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

ok no problem, but i hope it is clear why i am replacing x by \[\frac{b}{2}\]

OpenStudy (anonymous):

can you explain why you are replacing it?

OpenStudy (anonymous):

\[-(\frac{b}{2})^2+b\times \frac{b}{2}-75=25\] \[-(\frac{b}{2})^2+b\times \frac{b}{2}=100\] \[-\frac{b^2}{4}+\frac{b^2}{2}=100\] \[\frac{b^2}{4}=100\] \[b^2=400\] \[b=\pm\sqrt{400}\] \[b=\pm20\]

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

if you have a quadratic that looks like \[y=ax^2+bx+c\] then the maximum or minimum is always at the vertex (that is what it mean) and the first coordinate of the vertex is always \[x=-\frac{b}{2a}\] so i know that the maximum occurs when \[x=\frac{b}{2}\] in your example since a = -1. then i plug it in, and set it equal to 25 because you are told that the maximum is 25